OpenStudy (anonymous):

Fool's problem of the day, How many ordered pair of \((x,y)\) are there such that \(x, y \in \mathbb{Z} \) and \( \frac 2 x - \frac 3 y = \frac 1 4 \) Good luck!

Parth (parthkohli):

Oops, maybe this involves some trial and error :P

Parth (parthkohli):

What is \(\mathbb{Z}\) in here?

OpenStudy (anonymous):

http://en.wikipedia.org/wiki/Integer

Parth (parthkohli):

This works with -8 and -6

Parth (parthkohli):

If b is -8 then a will be -16

Parth (parthkohli):

I see a lot of solutions to this :/

OpenStudy (anonymous):

Not many, less than 25 I believe ;)

Parth (parthkohli):

24 or 23 or 22.. I guess

Parth (parthkohli):

The minimum of b is I guess -36 maximum is 84

OpenStudy (anonymous):

Don't guess, try to form a analytic approach :)

Parth (parthkohli):

Oh wait, it goes further.

Parth (parthkohli):

Your problem is hard.

OpenStudy (anonymous):

No spoiler please :)

Parth (parthkohli):

How do we do it by the way?

OpenStudy (anonymous):

It's problem of the day, it's supposed to be interesting. However my approach takes less than 10 seconds ;) I will post it later though.

OpenStudy (anonymous):

is the answer 5? @FoolForMath ?

Parth (parthkohli):

No Arnab.

OpenStudy (anonymous):

okay, we can get after simplification.. x=8y/(y+12) as x is an integer, 8y must be a multiple of y+12 8y=2*2*2*y so, here are possible equations y+12=1 y+12=2 y+12=4 y+12=8 y+12=2y y+12=4y y+12=8y as y +12 is a factor of 8y all the equations except the last one will satisfy for an integral value of y so, the answer is.. 6 ordered pairs

OpenStudy (anonymous):

not 5^^

OpenStudy (anonymous):

sorry, there will be other options too if RHS is negative y+12=-1 y+12=-2 ................. and so on.. for y to be an integer, there are another 5 solutions so, total 11

OpenStudy (anonymous):

8?

OpenStudy (anonymous):

is it right?

Parth (parthkohli):

No, not right

OpenStudy (anonymous):

the equation is equivalent to (x-8)*(y+12)=-96 so the answer is 11

Parth (parthkohli):

Na na

OpenStudy (anonymous):

is all the way wrong or just the 11?

OpenStudy (anonymous):

22

OpenStudy (anonymous):

I m pretty sur it's 22

OpenStudy (anonymous):

24

OpenStudy (anonymous):

is 24 correct answer.?

Parth (parthkohli):

No, you are close by the way

OpenStudy (anonymous):

I was able to find 11 ordered pairs that are solutions: (7, 84) (6,36) (5,20) (4,12) (2,4) (-4,-4) (-8,-6) (-16,-8) (-24,-9) (-40,-10) (-88,-11) I think there might be more, but im missing an idea.

OpenStudy (anonymous):

HINT: There are 23 of them.

OpenStudy (anonymous):

Real challenge is to find a quick analytic approach :)

OpenStudy (anonymous):

yeah i just found the other 12, making 23.

OpenStudy (anonymous):

Most people probably got that:\[8y-12x-xy = 0\]There is a way you can factor this to get an idea of how many solutions there are...

OpenStudy (anonymous):

That's right joe! Congrats man!

OpenStudy (anonymous):

ty :)